OpenStudy (anonymous):

Can someone explain to me what linear functions are i have having issues with understanding it, like system of linear equations. I need someone to break it down for me.

OpenStudy (anonymous):

linear functions usually given as \[f(x)=mx+b\] the fancy term would be a polynomial of degree 1

OpenStudy (anonymous):

called "linear" because if you graph \[y=mx+b\] you get a line

OpenStudy (anonymous):

\[m\] would be the slope, meaning how many units up or down you go in the y - direction for every one unit to the right in the x direction

OpenStudy (anonymous):

and \[b\] is where the graph crosses the y - axis because if you replace x by 0 you get \[f(0)=m\times 0+b=b\]meaning (0,b) is on the graph